OpenStudy (fantageplayer):

ok idk what zeros are A polynomial has been factored, as shown below: f(x) = (x − 3)(x + 8)(x − 11) What are the zeros of the polynomial? 3, 8, 11 3, −8, 11 −3, 8, −11 −3, 3, −8, 8, −11, 11

OpenStudy (fantageplayer):

@math&ing001

OpenStudy (math&ing001):

Let P(x) be a polynomial P(x) = (x-a)(x-b) a and b are the zeros

OpenStudy (ineedhelplz):

Split the function into three pieces: \[(x-3) = 0\] \[(x+8) = 0\] \[(x-11) = 0\] Solve for x

OpenStudy (fantageplayer):

so -3 8 and -11?

OpenStudy (ineedhelplz):

No

OpenStudy (math&ing001):

See what @Ineedhelplz did :)

OpenStudy (ineedhelplz):

3, 11, -8

OpenStudy (fantageplayer):

can u medal @Ineedhelplz

OpenStudy (fantageplayer):

wait what

OpenStudy (fantageplayer):

why did the signs switch?

OpenStudy (math&ing001):

You always take the opposite sign, it's like you're solving for x when P(x)=0

OpenStudy (fantageplayer):

ohhhhhhhhhhh

OpenStudy (fantageplayer):

thanks

OpenStudy (math&ing001):

np